# Robustly optimal dynamics for active matter reservoir computing

### v2

@@ Abstract
Information processing abilities of active matter are studied in the reservoir computing (RC) paradigm to infer the future state of a chaotic signal. We uncover an exceptional regime of agent dynamics that has been overlooked previously. It appears robustly optimal for performance under many conditions, thus providing valuable insights into computation with physical systems more generally. The key to forming effective mechanisms for information processing appears in the system’s intrinsic relaxation abilities. These are probed without actually enforcing a specific inference goal. The dynamical regime that achieves optimal computation is located just below a critical damping threshold, involving a relaxation with multiple stages, and is readable at the single-particle level. At the many-body level, it yields substrates robustly optimal for RC across varying physical parameters and inference tasks. A system in this regime exhibits a strong diversity of dynamic mechanisms under highly fluctuating driving forces. Correlations of agent dynamics can express a tight relationship between the responding system and the fluctuating forces driving it. As this model is interpretable in physical terms, it facilitates re-framing inquiries regarding learning and unconventional computing with a fresh rationale for many-body physics out of equilibrium.
@@ I Introduction
Computation is not limited to the digital domain: In materio computation is about (analogue) physical computing directly with matter, on a continuous domain of space and time [1, 2, 3]. Unlike Turing machines, a computation will not require sequential transitions between well-defined discrete internal states [4, 1, 5]. This is part of a wider push towards unconventional computing: using the intrinsic dynamics of physical systems for a computation can lessen the computational overhead arising from traditional electronic and digital systems [2, 6, 7, 8]. The idea is to to take inspiration from how nature computes [9], for example the brain [10, 11], import basic physical concepts and systems [1, 12, 13], and depart from the von Neumann computing paradigm, where information processing and memory storage are executed in separate compartments [14]. Hard condensed matter, i.e., magnetic systems, has been proven to offer high functionality for the computational tasks of prediction and pattern recognition: Topologically protected phases (skyrmions) are exploitable [15, 16, 17, 18, 19].
In the same spirit, can soft condensed matter present a complementary class of useful systems for computation and inference? Notably, soft materials are being used for sensing and actuating in robots [20]. Active matter systems [21] are interesting candidates because they are inherently out of equilibrium [22], rendering rich collective behavior at the macroscopic scale like flocking [23, 24]. How can we generally assess the suitability of soft or any kind of matter for computation?
@@
The reservoir computing (RC) paradigm is ideal for exploring these basic questions, as it exploits directly dynamical systems for information processing and inference (in real-time) [25, 26, 4]. The power of RC for scientific inquiry into physical computation is its generic and straightforward setup: Information of a low-dimensional input (e.g. a time series) is injected into the reservoir computing substrate, which is a high-dimensional nonlinear dynamical system that enters different dynamical states depending on the input signal and its history [27]. In the RC paradigm, the reservoir substrate is optimized based on heuristic global properties (such as the spectral radius in Echo State Networks). In the canonical case of utilizing a digital neural network, substrates are simply initialized in a random way. The readout layer [28] typically consists of a linear model trained using Ridge regression. This separation between training the nonlinear substrate and a simple readout layer constitutes the original discovery of RC as a method [25, 26, 4], avoiding costly backpropagation through time [28].
@@
RC renders state-of-the-art performance for time-series forecasting [29, 30, 31, 32], in particular, chaotic time series prediction [30]. Despite the plethora of other machine learning techniques available for time-series forecasting (nonlinear vector autoregression (“next-generation RC”) [33], NBEATS [34], NHiTS [35], transformer models [36] or LSTMs [37]), the highly generic method of RC using e.g. Echo State Networks has often remained an architecture of choice. It requires far less training data [31, 32] and allows for predictions several characteristic times of the chaotic system (Lyapunov times) ahead, in a closed-loop autoregressive fashion [31]. It is considered “model-free” in the sense that it can estimate properties of the input without the need for an explicit model thereof [38]. It can be implemented using generic substrates, e.g., enabling highly promising photonics-based machine learning [39, 40, 41, 42, 43].
@@
In RC, only a few general properties of the substrate are believed to play a role for maximum performance: the degree of nonlinearity, reproducibility (where similar inputs yield similar outputs), separability (where the reservoir differentiates distinct inputs into different outputs), and the echo state property (which refers to fading memory, meaning the current reservoir states depend only on recent history and not on initial conditions) [44, 4, 45, 46, 28, 47, 27].
However, many questions remain unresolved: For example, what is the optimal mix of nonlinearity and linearity in responses? [48] Do memory capacity and nonlinearity present trade-offs? [49, 48] What is the role of the time delay in response? [50, 51, 41, 40] How can input information be effectively injected into [33] and read out from [52] the system for more efficient RC? And how can interpretable frameworks be developed to integrate theory with practice and enhance the prevalence of RC? [53, 33].
@@
Yet, for reservoir computation in materio, deeper physical concepts remain hidden beneath the surface. For example, how do nonequilibrium phase diagrams and dissipative as well as correlative behavior relate to computational properties? Understanding this fundamental connection would be valuable for recognizing the true feasibility of general neuromorphic computing systems [19, 54, 12], and connect different realms of matter-based and digital computing, both conceptually and practically [55]. Artificial neural networks are usually employed as the model of RC, to develop mathematical theory; they are not particularly interpretable in these terms, essentially functioning as black-boxes [56]. Thus, a different paradigm for thinking about computation and learning is required when dealing with physical substrates [1, 13].
@@
RC has been demonstrated for soft matter, though extremely rarely, and developments remain in an early stage: Reported were cases of gel materials, soft macroscopic bodies, and biological cell cultures [57, 58, 59]. In one exotic example, water waves in a bucket were used as a proof-of-principle [60]. The usage of active matter RC is in its infancy [61]. Lymburn et al. simulated biologically inspired swarms avoiding an externally controlled predator as a reservoir computer [62], which motivated our work. Note that the converse idea of using a reservoir computer to control a swarm has also been investigated [63], and there is a growing interest in controlling active particles using artificial intelligence [64, 65, 66]. Wang et al. presented an experimental realization of RC with an array of active gold nano-particles driven by a laser [67]. Strong et al. designed RC around active hydrogen ions residing in an electroactive polymer gel, which can be excited by applying an electric field [68]. Two more experimental setups have been conceptually proposed by Jeggle and Wittkowski that leverage driving the active matter agents via spatiotemporally modulated fields [61]: micro-particles in a medium driven by ultrasonic pressure waves, and micro-particles with a symmetry-broken refractive index profile driven by light fields.
@@ II.2 Injecting information of a chaotic time-series into the system via external driving
is induced by the driver and experienced by an agent $i$. A cut-off radius $r_{d}$ is in place modeled by the Heaviside step function ($\theta(r_{d}-\left\|\mathbf{x}_{i}-\mathbf{x}_{d}\right\|)=1$ if $\left\|\mathbf{x}_{i}-\mathbf{x}_{d}\right\|\leq r_{d}$, and 00 otherwise). This term is added to the total force sum in Eq. 5 with a term $K_{d}\mathbf{F}_{d,i}$ that enters the $\tanh$ function in Eq. 6. The driver particle itself does not experience a force from the agents; its trajectory is fully determined by an external input time series. This time series is to be processed by the swarm reservoir computer to predict its future state.
@@ II.3 Extracting a set of coarse-grained reservoir state variables with Gaussian observation kernels
The closer agents are to a kernel’s center point, the higher their relative contribution to the kernel sum. At the beginning of a simulation, we place $M$ Gaussian observation kernels randomly on the square simulation canvas with lengths $l_{\text{box}}$. This is done by drawing $\mathbf{c}_{m}$ randomly from a uniform distribution $[0;l_{\text{box}})$, and $w_{m}$ from a Gaussian distribution with mean $\mu=0.0$ and standard deviation $\sigma=1/\sqrt{2}$. We note that in Ref. [62] a kernel placement protocol was employed that takes into account agent locations, but our simpler choice leads to similar results as verified in Fig. 12. For each time step $t$ of $T$ total simulation time steps we collect $3M$ observations, where $M$ is the number of kernels placed on the canvas, and save them in the observation matrix $\mathbf{X}\in\mathbb{R}^{3M\times T}$.
@@ II.5 Observables for driven nonequilibrium soft matter systems
where $\theta$ is the Heaviside step function ($\theta\left(r-r_{ij}\right)=1$ if $r_{ij}\leq r$, and 00 otherwise).
@@
$\Delta r_{c}(t)=\left\langle\frac{||\mathbf{x}(t)-\mathbf{x}_{\text{center}}||}{||\mathbf{x}(t_{0})||}\right\rangle_{N}$ (24)
@@
$\Delta v_{s}(t)=\left\langle\left(\frac{||\mathbf{v}(t)||-s}{s}\right)^{2}\right\rangle_{N}\,,$ (25)
which is given as the deviation of current agent speeds $||\mathbf{v}(t)||$ relative to the target agent speed $s$, averaged over all $N$ agents.
@@ II.6 Simulation and reservoir computing set-up
To perform reservoir computing, we simulate our active matter system in two runs: training and testing. For each run we choose different initial conditions for the chaotic input time series, leading to exponentially different trajectories due to their deterministically chaotic nature. Each run consists of an initial burn-in (equilibration) time of 1,000 steps that are neither recorded nor considered for the readout layer training and 50,000 recorded steps for the main simulation. After training the readout layer using the Gaussian kernel observations as described in Sect. II.3 and Sect. II.4, we predict $\Delta t_{\text{pred}}=25$ fixed integration time steps ahead by default at each time point in a non-autoregressive fashion, which corresponds to $\approx 0.45283$ Lyapunov times for the Lorenz attractor. The predicted Lyapunov time ahead for the other chaotic attractors is equivalent to that of the Lorenz-63 system.
@@ III.1 Around the critical damping regime: optimal performance
Figure 2: A critically damped dynamical regime shows optimal predictive performance for active matter reservoir computing. (a) Varying the parameters of the speed-controlling force reveals regimes with stark differences in predictive performance, corresponding to different agent dynamics. The heatmap shows varied target agent speeds $s\in[10^{-5};10^{2}]$ that a particle will accelerate or decelerate towards (see Eq. 4), and the strength coefficient $K_{sc}\in[10^{-5};10^{2}]$ of this force. In the center diagonal enclosed by the dashed yellow lines, there is a critically damped regime with suppressed agent excitations and a spatially well-confined collective swarm (circle, triangle, square symbols). Moving towards lower speed-controller strengths and higher target agent speeds yields a response that appears much more disordered than coherent-collective (diamond, star symbols). Higher speed-controller strengths and lower target agent speeds lead to arrested particle motion (nabla symbol), which becomes overdamped when choosing a smaller integration time step (see Fig. 16). The hexagon/“L” denotes the optimal dynamical regime presented in Ref. [62] (see Supplementary Video Tab. 1/2 for a snapshot and visualization). (b) Snapshots of swarms in different dynamical regimes marked as symbols in sub-figure (a); color indicates agent speed. Refer to Tab. 3 for corresponding videos. (c) Predictive performances $P$ are higher in the critically damped regime (triangle symbol, $P=0.85$) than in the underdamped regime (cross symbol, $P=0.52$). $P$ is defined as the correlation coefficient specified in Eq. 13 between the actual and the predicted time series. The gray curve shows the actual time series, while the colored curves show predictions $\Delta t=0.5$ time units ahead made with the active matter reservoir at each time point. (d) Agent-averaged mean squared displacements (MSDs) after the Lyapunov time of the Lorenz-63 attractor $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [100] for different speed-controller parameter combinations. Low MSDs separate the underdamped from the near-critical (and arrested) regime, and correlate with high predictive performances. They are thus reliable proxies for consistent, controlled agent responses to changing external stimuli. Parameter values $(K_{sc},s)$ for symbols: cross: (0.00005, 18.32981); diamond: (0.00379, 0.26367); square: (0.00886, 0.11288); triangle: (0.02069, 0.04833); circle: (0.04833, 0.02069); nabla: (0.26367, 0.00379).
@@
So far, it has been assumed that active matter reservoirs show optimal performance in a “critical” regime, around a fluid-to-gas-like phase transition [62], supporting the hypothesis of information processing being optimal around a phase transition [101]. Here we present a novel dynamical regime for improved reservoir computing with active matter: the critically damped regime. We uncover this regime by varying the parameters of the speed-controlling force $\mathbf{F}_{sc}$: the target agent speed $s$ and the strength $K_{sc}$ (see Eq. 4). Fig. 2a shows how well different active matter systems at different $(K_{sc},s)$ parameter combinations perform at predicting a chaotic input time series in a reservoir computing framework. We observe a broad region with high predictive performances which we define with yellow dashed lines ($s=r(K_{sc})^{m}$ with power $m=1$ and coefficients $r_{AB}\approx 0.2$ and $r_{BC}\approx 70$). The optimal ratio $r$ that delivers performances around $P=0.884$ is located at $K_{sc}=10^{-5}$ and $s=5.0\cdot 10^{-5}$. This is a significant increase in performance from the previously reported optimal regime [62] (reproduced as $P=0.719$) indicated by the hexagon/“L” symbol in Fig. 2a. Videos of the driven active matter systems (see Appendix E, Tab. 3) reveal that, upon driver excitation, particles do not embark on long trajectories or oscillate across the simulation box before returning to their steady-state positions. Instead, the speed-controller setting ensures that agents swiftly return to the box center. This is why we call this regime “critically damped”. The quick return and the higher order in this regime can also be observed from the snapshots in Fig. 2b (square, triangle, circle symbols). The swarm maintains a coherent droplet shape with low mean speeds (see Supplementary Fig. 19), even if perturbed by the driver.
@@
For $s\gg K_{sc}$, in the upper-left corner of Fig. 2a, agent speeds are only weakly regulated by the speed-controlling force. The overall response is much more disordered at the local particle level, which can be seen from the snapshot in Fig. 2a (cross symbol). Because agent momenta do not relax quickly, we call this regime underdamped. This notion becomes even more evident in the corresponding supplementary video. The predictive performance is on the order of $P=0.60$ and lower. Fig. 2c compares the strong differences in predictive performance between the actual and predicted time series in the near-critically damped ($P=0.85$) and the underdamped ($P=0.52$) regime.

### v3

@@ Abstract
Information processing abilities of active matter are studied in the reservoir computing (RC) paradigm to infer the future state of a chaotic signal. We uncover an exceptional regime of agent dynamics that has been overlooked previously. It appears robustly optimal for performance under many conditions, thus providing valuable insights into collective and physical computation more generally. The key to forming effective mechanisms for information processing appears in the system’s intrinsic relaxation abilities. These are probed without actually enforcing a specific inference goal. The dynamical regime that achieves optimal computation is located just below a critical damping threshold, involving a relaxation with multiple stages, and is readable at the single-particle level. At the many-body level, it yields substrates robustly optimal for RC across varying physical parameters and inference tasks. A system in this regime exhibits a strong diversity of dynamic mechanisms under highly fluctuating driving forces. Correlations of agent dynamics can express a tight relationship between the responding system and the fluctuating forces driving it. As this model is interpretable in physical terms, it facilitates re-framing inquiries regarding learning and computing with a fresh rationale for many-body physics far from equilibrium.
@@ I Introduction
Computation is not limited to the digital domain: In materio computation is about analog physical computing directly with matter, on a continuous domain of space and time [1, 2, 3]. Unlike Turing machines, a computation will not require sequential transitions between well-defined discrete internal states [4, 1, 5]. This is part of a wider push towards unconventional computing: using the intrinsic dynamics of physical systems for computation can lessen the computational overhead and resulting energy expenditure arising from traditional electronic and digital means [2, 6, 7, 8]. The idea is to to take inspiration from how nature computes [9], for example the brain [10, 11], import basic physical concepts and systems [1, 12, 13], and depart from the von Neumann computing paradigm, where information processing and memory storage are executed in separate compartments [14].
Can soft condensed matter or biological systems present a complementary class of systems adept for computation and inference? Does it rely on collectivity? Novel intelligent systems can hinge on, e.g., swarm intelligence to make composites of many, even heterogeneous parts act autonomously together – which is gaining rapid interest for artificial intelligence algorithms [15, 16] as well as robotics [17]. From a physical point of view, active matter systems [18] are interesting candidates because they are inherently out of equilibrium [19], rendering rich collective behavior at the macroscopic scale like swarming and flocking [20, 21]. Also, soft materials offer a way to distribute forces and integrate responses important for modern robotic sensing and actuation [22]. How can we generally assess the suitability of bio-inspired systems for both analog computation and learning algorithms, which represent significantly distinct contexts?
@@
The reservoir computing (RC) paradigm is ideal for exploring these basic questions. It directly exploits dynamical systems for information processing and inference in real-time [23, 24, 4]. The power of RC for scientific inquiry into physical computation is its generic and straightforward setup: Information of a low-dimensional input (e.g. a time series) is injected into the reservoir computing substrate, which is a high-dimensional nonlinear dynamical system that enters different dynamical states depending on the input signal and its history [25]. In the RC paradigm, the reservoir substrate is optimized based on heuristic global properties (such as the spectral radius in Echo State Networks). In the canonical case of utilizing a digital neural network, substrates are simply initialized in a random way. The readout layer [26] typically consists of a linear model trained using Ridge regression. This separation between training the nonlinear substrate and a simple readout layer constitutes the original discovery of RC as a method [23, 24, 4], avoiding costly backpropagation through time [26].
@@
Currently, RC renders state-of-the-art performance for time-series forecasting [27, 28, 29, 30], in particular, chaotic time series prediction [28]. Despite the plethora of other machine learning techniques available for time-series forecasting (nonlinear vector autoregression (“next-generation RC”) [31], NBEATS [32], NHiTS [33], transformer models [34] or LSTMs [35]), the highly generic method of RC using e.g. Echo State Networks has often remained an architecture of choice. It requires far less training data [29, 30] and allows for predictions several characteristic times of the chaotic system (Lyapunov times) ahead, in a closed-loop autoregressive fashion [29]. It is considered “model-free” in the sense that it can estimate properties of the input without the need for an explicit model thereof [36]. It can be implemented using generic substrates, e.g., enabling highly promising photonics-based machine learning [37, 38, 39, 40, 41].
@@
In RC, only a few general properties of the substrate are believed to play a role for maximum performance: the degree of nonlinearity, reproducibility (where similar inputs yield similar outputs), separability (where the reservoir differentiates distinct inputs into different outputs), and the echo state property (which refers to fading memory, meaning the current reservoir states depend only on recent history and not on initial conditions) [42, 4, 43, 44, 26, 45, 25]. However, many issues remain open: For example, the optimal mix of nonlinearity and linearity in responses [46]; if memory capacity and nonlinearity present trade-offs [47, 46]; time delay in response [48, 49, 39, 38]; and how information should be injected into [31] and read out from [50] the reservoir.
@@
Yet, for reservoir computation in materio or using biological-like swarms, deep physical concepts remain hidden beneath the surface. For example, how do nonequilibrium phase diagrams and dissipative as well as correlative dynamics relate to computational properties? Understanding these links would be valuable for recognizing the true feasibility of neuromorphic computing systems [51, 52, 12], and connecting different realms of matter-based and digital computing, both conceptually and practically [53]. Generally, interpretable RC frameworks should lead to better integration of theory with practice [54, 31]. Artificial neural networks have been employed as the standard substrate for RC; they are essentially black boxes [55]. A different paradigm for thinking about computation and learning is required when dealing with physical substrates [1, 13]. Notably, hard condensed matter, e.g., magnetic systems, has already offered high functionality for RC prediction and pattern recognition by exploiting topologically protected phases (skyrmions) [56, 57, 58, 59, 51].
Regarding living systems, it may not be clear whether the power of collective information processing is attributable to physical dynamics alone, or whether other factors may play a role [60, 61, 62, 63]. A basic physical model system for biologically-inspired RC can be enormously clarifying in these respects. Moreover, there is a rapidly growing interest in the use of biological matter for artificial intelligence, particularly within the RC paradigm, using self-organized brain organoids and objects alike [64, 65, 66].
@@
RC has hence been demonstrated for soft matter, though extremely rarely, and developments remain in an early stage: Reported were cases of gel materials, soft macroscopic bodies, and other examples of cell cultures [67, 68, 69]. In one exotic case, water waves in a bucket were used as a proof-of-principle [70]. The usage of active matter models for RC is in its infancy [71]. Lymburn et al. simulated biologically inspired swarms that avoid an externally controlled predator as a reservoir computer [72], which motivated our work. Note that the converse idea of using a reservoir computer to control a swarm has also been investigated [73], and there is a growing interest in controlling active particles using artificial intelligence [74, 75, 76]. Wang et al. presented an experimental realization of RC with an array of active gold nano-particles driven by a laser [77]. Strong et al. designed RC around active hydrogen ions residing in an electroactive polymer gel, which can be excited by applying an electric field [78]. Two more experimental setups have been conceptually proposed by Jeggle and Wittkowski that leverage driving the active matter agents via spatiotemporally modulated fields [71]: micro-particles in a medium driven by ultrasonic pressure waves, and micro-particles with a symmetry-broken refractive index profile driven by light fields.
@@ II.2 Injecting information of a chaotic time-series into the system via external driving
is induced by the driver and experienced by an agent $i$. A cut-off radius $r_{d}$ is in place modeled by the Heaviside step function ($\theta(r_{d}-\left\|\mathbf{x}_{i}-\mathbf{x}_{d}\right\|)=1$ if $\left\|\mathbf{x}_{i}-\mathbf{x}_{d}\right\|\leq r_{d}$, and 0 otherwise). This term is added to the total force sum in Eq. 5 with a term $K_{d}\mathbf{F}_{d,i}$ that enters the $\tanh$ function in Eq. 6. The driver particle itself does not experience a force from the agents; its trajectory is fully determined by an external input time series. This time series is to be processed by the swarm reservoir computer to predict its future state.
@@ II.3 Extracting a set of coarse-grained reservoir state variables with Gaussian observation kernels
The closer agents are to a kernel’s center point, the higher their relative contribution to the kernel sum. At the beginning of a simulation, we place $M$ Gaussian observation kernels randomly on the square simulation canvas with lengths $l_{\text{box}}$. This is done by drawing $\mathbf{c}_{m}$ randomly from a uniform distribution $0;l_{\text{box}})$, and $w_{m}$ from a Gaussian distribution with mean $\mu=0.0$ and standard deviation $\sigma=1/\sqrt{2}$. We note that in Ref. [72] a kernel placement protocol was employed that takes into account agent locations, but our simpler choice leads to similar results as verified in Fig. 12. For each time step $t$ of $T$ total simulation time steps, we collect $3M$ observations, where $M$ is the number of kernels placed on the canvas, and save them in the observation matrix $\mathbf{X}\in\mathbb{R}^{3M\times T}$.
@@ II.5 Observables for driven nonequilibrium soft matter systems
where $\theta$ is the Heaviside step function ($\theta\left(r-r_{ij}\right)=1$ if $r_{ij}\leq r$, and 0 otherwise).
@@
[…]
@@
[…]
which is given as the deviation of current agent speeds $\mathbf{v}(t)||$ relative to the target agent speed $s$, averaged over all $N$ agents.
@@ II.6 Simulation and reservoir computing set-up
To perform reservoir computing, we simulate our active matter system in two runs: training and testing. For each run, we choose different initial conditions for the chaotic input time series, leading to exponentially different trajectories due to their deterministically chaotic nature. Each run consists of an initial burn-in (equilibration) time of 1,000 steps that are neither recorded nor considered for the readout layer training and 50,000 recorded steps for the main simulation. After training the readout layer using the Gaussian kernel observations as described in Sect. II.3 and Sect. II.4, we predict $\Delta t_{\text{pred}}=25$ fixed integration time steps ahead by default at each time point in a non-autoregressive fashion, which corresponds to $\approx 0.45283$ Lyapunov times for the Lorenz attractor. The predicted Lyapunov time ahead for the other chaotic attractors is equivalent to that of the Lorenz-63 system.
@@ III.1 Around the critical damping regime: optimal performance
Figure 2: A critically damped dynamical regime shows optimal predictive performance for active matter reservoir computing. (a) Varying the parameters of the speed-controlling force reveals regimes with stark differences in predictive performance, corresponding to different agent dynamics. The heatmap shows varied target agent speeds $s\in[10^{-5};10^{2}]$ that a particle will accelerate or decelerate towards (see Eq. 4), and the strength coefficient $K_{sc}\in[10^{-5};10^{2}]$ of this force. In the center diagonal enclosed by the dashed yellow lines, there is a critically damped regime with suppressed agent excitations and a spatially well-confined collective swarm (circle, pyramid, square symbols). Moving towards lower speed-controller strengths and higher target agent speeds yields a response that appears much more disordered than coherent-collective (diamond, star symbols). Higher speed-controller strengths and lower target agent speeds lead to arrested particle motion (nabla symbol), which becomes overdamped when choosing a smaller integration time step (see Fig. 16). The hexagon/“L” denotes the optimal dynamical regime presented in Ref. [72] (see Supplementary Video Tab. 1/2 for a snapshot and visualization). (b) Snapshots of swarms in different dynamical regimes marked as symbols in sub-figure (a); color indicates agent speed. Refer to Tab. 3 for corresponding videos. (c) Predictive performances $P$ are higher in the critically damped regime (pyramid symbol, $P=0.88$) than in the underdamped regime (cross symbol, $P=0.52$). $P$ is defined as the correlation coefficient specified in Eq. 13 between the actual and the predicted time series. The gray curve shows the actual time series, while the colored curves show predictions $\Delta t=0.5$ time units ahead made with the active matter reservoir at each time point. (d) Agent-averaged mean squared displacements (MSDs) after the Lyapunov time of the Lorenz-63 attractor $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [110] for different speed-controller parameter combinations. Low MSDs separate the underdamped from the near-critical (and arrested) regime, and correlate with high predictive performances. They are thus reliable proxies for consistent, controlled agent responses to changing external stimuli. Parameter values $(K_{sc},s)$ for symbols: cross: (0.00005, 18.32981); diamond: (0.00379, 0.26367); square: (0.00886, 0.11288); pyramid: (0.02069, 0.04833); circle: (0.04833, 0.02069); nabla: (0.26367, 0.00379).
@@
So far, it has been assumed that active matter reservoirs show optimal performance in a “critical” regime, around a fluid-to-gas-like phase transition [72], supporting the hypothesis of information processing being optimal around a phase transition [111]. Here we present a novel dynamical regime for improved reservoir computing with active matter: the critically damped regime. We uncover this regime by varying the parameters of the speed-controlling force $\mathbf{F}_{sc}$: the target agent speed $s$ and the strength $K_{sc}$ (see Eq. 4). Fig. 2a shows how well different active matter systems at different $(K_{sc},s)$ parameter combinations perform at predicting a chaotic input time series in a reservoir computing framework. We observe a broad region with high predictive performances which we define with yellow dashed lines ($s=r(K_{sc})^{m}$ with power $m=1$ and coefficients $r_{AB}\approx 0.2$ and $r_{BC}\approx 70$). The optimal ratio $r$ that delivers performances around $P=0.884$ is located at $K_{sc}=10^{-5}$ and $s=5.0\cdot 10^{-5}$. This is a significant increase in performance from the previously reported optimal regime [72] (reproduced as $P=0.719$) indicated by the hexagon/“L” symbol in Fig. 2a. Videos of the driven active matter systems (see Appendix E, Tab. 3) reveal that, upon driver excitation, particles do not embark on long trajectories or oscillate across the simulation box before returning to their steady-state positions. Instead, the speed-controller setting ensures that agents swiftly return to the box center. This is why we call this regime “critically damped”. The quick return and the higher order in this regime can also be observed from the snapshots in Fig. 2b (square, pyramid, circle symbols). The swarm maintains a coherent droplet shape with low mean speeds (see Supplementary Fig. 19), even if perturbed by the driver.
@@
For $s\gg K_{sc}$, in the upper-left corner of Fig. 2a, agent speeds are only weakly regulated by the speed-controlling force. The overall response is much more disordered at the local particle level, which can be seen from the snapshot in Fig. 2a (cross symbol). Because agent momenta do not relax quickly, we call this regime underdamped. This notion becomes even more evident in the corresponding supplementary video. The predictive performance is on the order of $P=0.60$ and lower. Fig. 2c compares the strong differences in predictive performance between the actual and predicted time series in the near-critically damped ($P=0.88$) and the underdamped ($P=0.52$) regime.